-2 3/4 - (-3 5/6) =
Found 2 solutions by Alan3354, math_helper:
Answer by Alan3354(69443)

It's arithmetic.
Make an effort.

Answer by math_helper(2461)

For problems like this, I like to write the mixed numbers as improper fractions, while remembering that numbers like -2 3/4 are (-1)(2 + 3/4):

-2 3/4 - (-3 5/6) = -11/4 + 23/6
and then you need to put them over a common denominator (in this case, 12):
= -33/12 + 46/12
= 13/12
= 1 1/12
or about 1.0833333